# Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models

### v1

@@ 4.2 Generation
4.3 Additional results with CIFAR10 and Tiny ImageNet
Since MNIST-like datasets are rather simple, it is difficult to assess the true gain in performance resulting from the introduction of ECCs proposed in our model. This section presents additional results using CIFAR10 and Tiny ImageNet, which contain colored images with more intricate shapes, patterns, and greater diversity than the previous datasets. We trained uncoded and coded models using different configurations to gain intuition regarding the effect of introducing the ECC. For a reference, in the case of the DVAE++ [30], the authors needed 128 binary latent variables to achieve state-of-the-art performance in generation and reconstruction for this dataset. They employed a more intricate model than the one introduced in this study, featuring Boltzmann Machine priors. In Fig. 5 we show examples of reconstruction using different configurations of the model and in Fig. 6 we show examples of randomly generated images. Additional results are provided in Appendices D and E. The results are consistent with those presented in previous sections; but in this case, the difference in performance is even more pronounced. We observe that the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space. Nevertheless, the coded DVAE shows particular promise for learning low-dimensional discrete latent representations in complex datasets. Note that we used a rather simple architecture as we want to focus on the gain obtained only by introducing ECCs in the latent space.
@@ 5 Beyond repetition codes
  • Differentiability. Both encoding and decoding processes must be differentiable w.r.t the inputs to enable gradient computation and backpropagation.
Figure 7: Graphical model of the hierarchical coded DVAE.
Since Shannon’s seminal work [25], researchers have developed effective ECC schemes that meet these properties, including state-of-the-art ECCs such as Low Density Parity Check (LDPC) codes [8], or polar codes [2]. Efforts have also focused on developing efficient SISO decoders, such as the sum-product algorithm [15].
Inspired by polar codes [2], we present a hierarchical coded DVAE with two layers of latent bits. In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$. Simultaneously, the vector $\boldsymbol{m}_{2}$ is linearly combined with $\boldsymbol{m}_{1}$ using modulo 2 operations $\big{(}\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}\big{)}$ and then encoded using a repetition code, yielding $\boldsymbol{c}_{2}$ and $\boldsymbol{z}_{2}$. Both soft vectors are concatenated and fed to the decoder NN to generate $\boldsymbol{x}$. The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model. Inference follows a similar approach to the coded DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch. This hierarchical structure allows the model to effectively separate high-level information from finer details, as we show in the results presented in Appendix G.
@@ 6 Conclusion
This paper presents the first proof-of-concept demonstration that safeguarding latent information with ECCs within deep generative models holds promise for enhancing overall performance. By integrating redundancy into the latent space, the variational family can effectively refine the inference network’s output according to the structure of the ECC. Our findings underscore the efficacy of simple and efficient ECCs, like repetition codes, showcasing remarkable improvements over a lightweight version of the DVAE introduced in [30].
@@
Furthermore, our work reveals numerous avenues for future research. Firstly, investigating decoder architectures capable of efficiently utilizing the correlations and structure introduced by the ECCs, in contrast to the feed-forward networks employed in this study. We also contemplate exploring more complex and robust coding schemes, conducting theoretical analyses aligned with Shannon’s channel capacity and mutual information concepts to determine the fundamental parameters of the ECC needed to achieve reliable variational inference, exploring different modulations, and integrating these concepts into state-of-the-art models based on discrete representations.
7 Acknowledgements
@@
The work was partially supported by the Ministerio de Ciencia, Innovación y Universidades, Agencia Estatal de Investigación, under grant PID2021-123182OB-I00, MCIN/AEI/10.13039/501100011033, co-financed by European Regional Development Fund (ERDF), “A way of making Europe”. Pablo M. Olmos also acknowledges the support by BBVA Foundation (Beca Leonardo a Investigadores y Creadores Culturales 2024) and by the Comunidad de Madrid under grants IND2022/TIC-23550 and ELLIS Unit Madrid. David Mitchell acknowledges support from the National Science Foundation under Grant No. CCF-2145917 and travel support from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 951847. Grace Villacrés acknowledges support from Comunidad de Madrid within the 2023-2026 agreement with Universidad Rey Juan Carlos for the granting of direct subsidies for the promotion and encouragement of research and technology transfer, Line of Action A: Emerging Doctors, under Project OrdeNGN.
Appendix A Architecture
@@
In this section, we detail the architecture used to obtain the experimental results with FMNIST and MNIST (28x28 gray-scale images). Note that across experiments we only modify the output layer of the encoder and the input layer of the decoder to adapt to the different configurations of the model. This modification leads to a minimal alteration in the total number of parameters. In Section H.2, we conduct an ablation study on the number of trainable parameters to show that the enhancement in performance is not attributed to the increased dimensionality introduced by redundancy. For the additional CIFAR10 experiments, we change the input of the encoder and the output of the decoder to process the 32x32 color images. For the Tiny ImageNet experiments, we do the same to process 64x64 color images. The rest of the architecture remains unchanged.
A.1 Encoder
@@
The encoder NN consists of 3 convolutional layers followed by two fully connected layers. We employed Leaky ReLU as the intermediate activation function and a Sigmoid as the output activation, as the encoder outputs bit probabilities. The full architecture is detailed in Fig. 8.
@@
A.2 Decoder
The decoder architecture is inspired by the one proposed in [24]. It is composed of two fully connected layers, followed by transposed convolutional layers with residual connections and Squeeze-and-Excitation (SE) layers [10]. We employed Leaky ReLU as the intermediate activation function and a Sigmoid as output activation, given that we consider datasets with gray-scale images. The complete architecture is detailed in Fig. 9.
@@ A.2 Decoder
Appendix B FMNIST results
@@ B.1 Training
Across all cases, coded models achieve superior bounds. The main differences in the ELBO come from the different performances in reconstruction. As we have observed in the different experiments, coded models are capable of generating more detailed images and accurate reconstructions. Introducing the repetition code also leads to smaller KL values, indicating that the posterior latent features are disentangled and less correlated.
@@
B.2 Reconstruction and generation
In this section, we augment the results presented in the main text, including outcomes obtained with the training dataset in Table 1. We include again the results obtained in the test to facilitate comparison. The results remain consistent across the two data partitions, and the analysis conducted for the test set also applies to training data.
In all the cases, the coded models yield higher PSNR values than their uncoded counterparts, indicating a superior performance in reconstruction. We observe a general improvement in PSNR as we increase the number of information bits (i.e., as we augment the latent dimensionality of the model) and decrease the code rate (i.e., as we introduce more redundancy).
As we discussed in the main text, the PSNR does not account for the semantic errors committed by the model. Therefore, we additionally report the semantic accuracy and the confident semantic accuracy. While the reconstruction accuracy is computed across the entire dataset partitions, for the confident accuracy, we only consider those images projected into a latent vector with a probability exceeding $0.4$. We observe that coded models better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
In Fig 13 we include additional examples of randomly generated images using different model configurations. We observe that coded models can generate more detailed and diverse images than their uncoded counterparts.
@@ B.2 Reconstruction and generation
uncoded 5 14.490 0.541 0.541 14.477 0.536 0.536
coded 5/50 16.375 0.656 0.702 16.241 0.647 0.700
coded 5/80 16.824 0.694 0.751 16.624 0.688 0.748
coded 5/100 17.001 0.708 0.760 16.702 0.700 0.757
uncoded 8 15.644 0.601 0.602 15.598 0.594 0.595
coded 8/80 17.877 0.769 0.842 17.318 0.750 0.816
coded 8/160 18.828 0.807 0.878 17.713 0.783 0.831
coded 8/240 19.345 0.831 0.921 17.861 0.799 $\mathbf{0.893}$
uncoded 10 16.053 0.650 0.652 16.000 0.644 0.648
coded 10/100 18.827 0.813 0.885 17.694 0.790 0.850
coded 10/200 19.937 0.846 0.897 18.009 0.814 0.871
coded 10/300 $\mathbf{20.529}$ $\mathbf{0.855}$ $\mathbf{0.907}$ $\mathbf{18.111}$ $\mathbf{0.817}$ 0.870
Appendix C MNIST results
@@ C.2 Reconstruction
In Table 2 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets. In all the cases, coded models yield higher PSNR values, indicating a superior performance in reconstruction. This improvement is also evident through visual inspection of Fig. 17, where the coded models better capture the details in the images. As in FMNIST, we observe a general improvement of the PSNR as we increase the number of information bits and decrease the code rate.
As we discussed in the main text, the PSNR does not account for the semantic errors committed by the model. Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones. We also provide a confident reconstruction accuracy, for which we do not count errors when the MAP value of $q(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4. In light of the results, we argue that introducing ECCs in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, outperforming the uncoded models in all the cases.
@@
uncoded 5 13.483 0.702 0.703 13.483 0.701 0.702 0.277
coded 5/50 14.983 0.887 0.923 14.888 0.887 0.920 2.073
coded 5/80 15.436 0.899 0.936 15.263 0.895 0.929 2.237
coded 5/100 15.590 0.905 0.931 15.352 0.898 0.924 2.382
uncoded 8 14.530 0.860 0.864 14.490 0.860 0.868 0.513
coded 8/80 16.878 0.937 0.964 16.042 0.912 0.947 3.105
coded 8/160 18.108 0.957 0.974 16.497 0.927 0.951 3.645
coded 8/240 19.984 0.967 0.978 16.688 0.936 0.957 3.881
uncoded 10 14.879 0.888 0.891 14.816 0.887 0.890 0.636
coded 10/100 17.584 0.945 0.972 16.795 0.928 $\mathbf{0.968}$ 4.080
coded 10/200 20.060 0.973 0.977 16.863 0.932 0.944 4.411
coded 10/300 $\mathbf{21.083}$ $\mathbf{0.979}$ $\mathbf{0.984}$ $\mathbf{17.114}$ $\mathbf{0.941}$ 0.945 4.810
We also report the average entropy of the variational posterior over the test set in Table 2. If we analyze the entropy together with the semantic accuracy, we can argue that coded VAE is aware that multiple vectors might be related to the same image class, and that the posterior shows larger uncertainties for images for which the model has not properly identified the class. We illustrate this argument in Fig. 18, where we show some images selected so that the MAP latent word of $q(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors. We show the reconstruction of the 4 most probable latent vectors and their corresponding probabilities. Observe that the uncoded model is confident no matter the reconstruction outcome, while in the coded posterior, the uncertainty is much larger.
@@
C.3 Generation
In this section, we evaluate the model in the image generation task. In Fig. 19, we show examples of randomly generated images using different model configurations in MNIST. These results are consistent with the ones obtained in reconstruction, and with the ones obtained for FMNIST, as we observe that the coded models can generate more detailed and diverse images.
The improved inference provided by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational posterior distribution. Table 4 reports the BER and WER for MNIST. Remarkably, for the same number of latent bits, coded models reduce both the BER and WER w.r.t. the uncoded case. Note also that the error rates grow with the number of latent bits, but this is expected due to the increased complexity of the inference process.

### v2

@@ 5.2 Reconstruction
Coded models, in contrast, improve semantic accuracy and exhibit higher entropy. This indicates that (i) the Coded-DVAE recognizes multiple latent vectors that might be related to a given semantic class, and (ii) the model’s posterior demonstrates significant uncertainty (high entropy) for images where the class is not well identified. This is illustrated in Fig. 6, where we show images with class reconstruction errors caused by the MAP latent word from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The reconstructions of the three most probable latent vectors, along with their respective probabilities, are displayed. Notably, the uncoded model shows high confidence regardless of the reconstruction result, while the coded posterior shows greater uncertainty. Note that this does not imply an uninformative distribution, as information bits remain recoverable. Additionally, increasing the number of latent bits does not lead to an excessive increase in entropy, despite the exponential growth in the number of vectors.
5.3 Results on CIFAR10 and ImageNet
Since MNIST-like datasets are relatively simple, it is challenging to fully assess the benefits of introducing ECCs in our model. To address this, we present additional results on more complex datasets, CIFAR10 and Tiny ImageNet, which feature colored images with more intricate shapes, patterns, and greater diversity. We trained both uncoded and coded models with different configurations to better understand the impact of the ECC. For context, the DVAE++ model [Vahdat et al., 2018b] required 128 binary latent variables to achieve state-of-the-art performance on these datasets, using a more complex model with Boltzmann Machine priors. In Fig. 7, we show reconstruction examples, and in Fig. 8, we present randomly generated images. Additional results are available in Appendices F and G. The results align with those from previous sections, but the performance difference is even more pronounced in this case. We find that the uncoded DVAE struggles to decouple spatial information in the images, while the Coded-DVAE shows particular promise for learning low-dimensional discrete latent representations, even for complex datasets. It is important to note that we used a simple architecture to isolate the gains achieved purely through the introduction of ECCs in the latent space, keeping a simple independent prior.
@@ 6 Beyond repetition codes
We have presented compelling proof-of-concept results that incorporating ECCs, like repetition codes, into DVAEs can improve performance. We believe this opens a new path for designing latent probabilistic models with discrete latent variables. Although a detailed analysis of the joint design of ECC and encoder-decoder networks is beyond the scope of this work, in Appendix I we outline key properties that any ECCs must satisfy to be integrated within this framework.
Drawing inspiration from polar codes [Arikan, 2009], we introduce a hierarchical Coded-DVAE with two layers. The graphical model is shown in Fig. 9. In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$. Simultaneously, the vector $\boldsymbol{m}_{2}$ is linearly combined with $\boldsymbol{m}_{1}$ using modulo 2 operations $\big{(}\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}\big{)}$ and then encoded using a repetition code, yielding $\boldsymbol{c}_{2}$ and $\boldsymbol{z}_{2}$. Both soft vectors are concatenated and fed to the decoder to generate $\boldsymbol{x}$. Since $\boldsymbol{m}_{1}$ appears in both generative branches, it receives stronger protection. Inference follows a similar approach to the Coded-DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch. This hierarchical structure allows the model to separate high-level information from finer details, as we show in the results presented in Appendix J. The bits in $\boldsymbol{m}_{1}$ offer stronger protection, allowing the model to encode the semantic information (type of clothing in FMNIST). As a result, during inference, it is more likely to project the image into a latent vector $\boldsymbol{m}_{1}$ that represents the correct semantic class. In contrast, $\boldsymbol{m}_{2}$ provides weaker protection and encodes image details that are less crucial for reconstruction. Fig. 9 demonstrates that by keeping $\boldsymbol{m}_{1}$ fixed (semantic information) and sampling $\boldsymbol{m}_{2}$ randomly (fine details) we can generate diverse images within the same semantic class.
Figure 9: Hierarchical Coded-DVAE. Graphical model (right) and generated images with a 5/100 code per branch (left). Here, $\boldsymbol{m}_{1}$ is fixed, and $\boldsymbol{m}_{2}$ is randomly sampled.
@@ 7 Conclusion
This paper presents the first proof-of-concept demonstration that safeguarding latent information with ECCs within deep generative models holds promise for enhancing overall performance. By integrating redundancy into the latent space, the variational family can effectively refine the inference network’s output according to the structure of the ECC. Our findings underscore the efficacy of simple and efficient ECCs, like repetition codes, showcasing remarkable improvements over a lightweight version of the DVAE introduced in Vahdat et al. [2018b].
@@
Furthermore, our work reveals numerous avenues for future research. Firstly, investigating architectures capable of efficiently utilizing the correlations and structure introduced by the ECCs, in contrast to the feed-forward networks employed in this study. We also contemplate exploring more complex and robust coding schemes, conducting theoretical analyses aligned with Shannon’s channel capacity and mutual information to determine the fundamental parameters of the ECC needed to achieve reliable inference, exploring different modulations, and integrating these concepts into state-of-the-art models based on discrete representations.
Acknowledgements.
@@
María Martínez-García was supported by the Generación de Conocimiento grant PID2022-142506NA-I00. Grace Villacrés was supported by the Comunidad de Madrid within the 2023-2026 agreement with Universidad Rey Juan Carlos for the granting of direct subsidies for the promotion, encouragement of research and technology transfer, line of Action A Emerging Doctors, under Project OrdeNGN (Ref. F1177). David Mitchell was supported by the National Science Foundation under Grant No. CCF-2145917 and he acknowledges travel support from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 951847. Pablo M. Olmos was supported by the Comunidad de Madrid IND2022/TIC-23550, IDEA-CM project (TEC-2024/COM-89), the ELLIS Unit Madrid (European Laboratory for Learning and Intelligent Systems), the 2024 Leonardo Grant for Scientific Research and Cultural Creation from the BBVA Foundation, and by projects MICIU/AEI/10.13039/501100011033/FEDER and UE (PID2021-123182OB-I00; EPiCENTER). We would like to thank Isabel Valera for her valuable feedback on earlier drafts of the manuscript, as well as for the discussions that helped us improve the presentation of our method to the machine learning community.
Appendix A Uncoded training algorithm
The following pseudo-code describes the training process for the uncoded DVAE. It’s important to note that the main difference from the training of the Coded-DVAE lies in the fact that the encoder directly outputs $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{m}|\boldsymbol{x}_{i})$, which is used to sample $\boldsymbol{z}$. Therefore, we skip the soft decoding and coding steps.
Algorithm 1 Training the model with uncoded inference.
1:  Input: training data $\boldsymbol{x}_{i}$.
2:  repeat
4:     $\boldsymbol{z}\leftarrow$ sample from (5)
6:     Compute ELBO according to (3)
7:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
8:  until convergence
Appendix B Coded training algorithm
The following pseudo-code describes the training process for the Coded-DVAE. Here, we utilize soft decoding to leverage the added redundancy and retrieve the marginal posteriors of the information bits $\boldsymbol{m}$, correcting potential errors in $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})$. We then apply the soft encoding technique to incorporate the structure of the code and sample $\boldsymbol{z}$ using the reparameterization trick as described in (5).
Algorithm 2 Training the Coded-DVAE with repetition codes.
1:  Input: training data $\boldsymbol{x}_{i}$, matrix $\mathbf{G}$.
2:  repeat
6:     $\boldsymbol{z}\leftarrow$ sample from (5)
8:     Compute ELBO according to (3)
9:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
10:  until convergence
Appendix C Architecture
@@
In this section, we detail the architecture used to obtain the experimental results with FMNIST and MNIST (28x28 gray-scale images). Note that, across the experiments, we only modify the output layer of the encoder and the input layer of the decoder to adapt to the different configurations of the model. This modification leads to a minimal alteration in the total number of parameters. In Section K.2, we conduct an ablation study on the number of trainable parameters to show that the enhancement in performance is not attributed to the increased dimensionality introduced by redundancy.
For the additional CIFAR10 experiments, we change the input of the encoder and the output of the decoder to process the 32x32 color images. For the Tiny ImageNet experiments, we do the same to process 64x64 color images. The rest of the architecture remains unchanged.
These architectures are comprehensively described in the following subsections.
C.1 Encoder
@@
The encoder NN consists of 3 convolutional layers followed by two fully connected layers. We employed Leaky ReLU as the intermediate activation function and a Sigmoid as the output activation, since the encoder outputs bit probabilities. The full architecture is detailed in Fig. 10.
@@
C.2 Decoder
The decoder architecture is inspired by the one proposed in Schuster and Krogh [2023]. It is composed of two fully connected layers, followed by transposed convolutional layers with residual connections and Squeeze-and-Excitation (SE) layers [Hu et al., 2018]. We employed Leaky ReLU as the intermediate activation function and a Sigmoid as output activation, given that we consider datasets with gray-scale images. The complete architecture is detailed in Fig. 11.
@@ C.2 Decoder
Appendix D FMNIST results
@@ D.1 Training
Across all cases, coded models achieve superior bounds. The main differences in the ELBO come from the different performances in reconstruction. As we have observed in the different experiments, coded models are capable of generating more detailed images and accurate reconstructions. Introducing the repetition code also leads to smaller KL values, indicating that the posterior latent features are potentially disentangled and less correlated.
@@
D.2 Reconstruction and generation
In this section, we augment the results presented in the main text, including outcomes obtained with the training dataset in Table 1. We include again the results obtained in the test to facilitate comparison. Additionally, we assess reconstruction quality using the Structural Similarity Index Measure (SSIM) to provide a more comprehensive evaluation. The results remain consistent across the two data partitions, and the analysis conducted for the test set also applies to training data.
In all the cases, the coded models yield higher PSNR and SSIM values than their uncoded counterparts, indicating a superior performance in reconstruction. We observe a general improvement in PSNR and SSIM as we increase the number of information bits (i.e., as we augment the latent dimensionality of the model) and decrease the code rate (i.e., as we introduce more redundancy).
As we discussed in the main text, the reconstruction metrics such as the PSNR and SSIM do not account for the semantic errors committed by the model. Therefore, we additionally report the semantic accuracy and the confident semantic accuracy. While the reconstruction accuracy is computed across the entire dataset partitions, for the confident accuracy, we only consider those images projected into a latent vector with a probability exceeding $0.4$. We observe that coded models better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
@@ D.2 Reconstruction and generation
In Fig. 15, we include additional examples of randomly generated images using different model configurations. We observe that coded models can generate more detailed and diverse images than their uncoded counterparts.
@@
uncoded 5 14.490 0.438 0.541 0.541 14.477 0.437 0.536 0.536
coded 5/50 16.375 0.559 0.656 0.702 16.241 0.551 0.647 0.700
coded 5/80 16.824 0.585 0.694 0.751 16.624 0.574 0.688 0.748
coded 5/100 17.001 0.596 0.708 0.760 16.702 0.580 0.700 0.757
uncoded 8 15.644 0.513 0.601 0.602 15.598 0.509 0.594 0.595
coded 8/80 17.877 0.647 0.769 0.842 17.318 0.619 0.750 0.816
coded 8/160 18.828 0.690 0.807 0.878 17.713 0.641 0.783 0.831
coded 8/240 19.345 0.713 0.831 0.921 17.861 0.653 0.799 $\mathbf{0.893}$
uncoded 10 16.053 0.542 0.650 0.652 16.000 0.536 0.644 0.648
coded 10/100 18.827 0.690 0.813 0.885 17.694 0.639 0.790 0.850
coded 10/200 19.937 0.735 0.846 0.897 18.009 0.659 0.814 0.871
coded 10/300 $\mathbf{20.529}$ $\mathbf{0.754}$ $\mathbf{0.855}$ $\mathbf{0.907}$ $\mathbf{18.111}$ $\mathbf{0.662}$ $\mathbf{0.817}$ 0.870
Appendix E MNIST results
@@ E.2 Reconstruction
In Table 2 we quantify the quality of the reconstructions measuring the PSNR and the SSIM in both training and test sets. In all the cases, coded models yield higher PSNR and SSIM values, indicating a superior performance in reconstruction. This improvement is also evident through visual inspection of Fig. 19, where the coded models better capture the details in the images. Similar to the results observed on FMNIST, we find that both PSNR and SSIM generally improve as the number of information bits increases and the code rate decreases.
@@
uncoded 5 13.491 0.434 0.702 0.703 13.483 0.430 0.701 0.702 0.277
coded 5/50 14.983 0.601 0.887 0.923 14.888 0.598 0.887 0.920 2.073
coded 5/80 15.436 0.641 0.899 0.936 15.263 0.634 0.895 0.929 2.237
coded 5/100 15.590 0.652 0.905 0.931 15.352 0.641 0.898 0.924 2.382
uncoded 8 14.530 0.558 0.860 0.864 14.490 0.558 0.860 0.868 0.513
coded 8/80 16.878 0.739 0.937 0.964 16.042 0.699 0.912 0.947 3.105
coded 8/160 18.108 0.802 0.957 0.974 16.497 0.736 0.927 0.951 3.645
coded 8/240 19.984 0.838 0.967 0.978 16.688 0.752 0.936 0.957 3.881
uncoded 10 14.879 0.591 0.888 0.891 14.816 0.589 0.887 0.890 0.636
coded 10/100 17.584 0.777 0.945 0.972 16.795 0.744 0.928 $\mathbf{0.968}$ 4.080
coded 10/200 20.060 0.875 0.973 0.977 16.863 0.765 0.932 0.944 4.411
coded 10/300 $\mathbf{21.083}$ $\mathbf{0.902}$ $\mathbf{0.979}$ $\mathbf{0.984}$ $\mathbf{17.114}$ $\mathbf{0.781}$ $\mathbf{0.941}$ 0.945 4.810
As we discussed in the main text, reconstruction metrics such as PSNR and SSIM do not account for the semantic errors committed by the model. Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones. We also provide a confident reconstruction accuracy, for which we do not count errors when the MAP value of $q(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4. In light of the results, we argue that introducing ECCs in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, outperforming the uncoded models in all the cases.
We also report the average entropy of the variational posterior over the test set in Table 2. If we consider the entropy together with the semantic accuracy, we can argue that coded VAE is aware that multiple vectors might be related to the same image class, and that the posterior shows larger uncertainties for images for which the model has not properly identified the class. We illustrate this argument in Fig. 20, where we show some images selected so that the MAP latent word of $q(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors. We show the reconstruction of the 4 most probable latent vectors and their corresponding probabilities. Observe that the uncoded model is confident no matter the reconstruction outcome, while in the coded posterior, the uncertainty is much larger.
@@
E.3 Generation
In this section, we evaluate the model in the image generation task. In Fig. 21, we show examples of randomly generated images using different model configurations in MNIST. These results are consistent with the ones obtained in reconstruction, and with the ones obtained for FMNIST, where we observe that the coded models can generate more detailed and diverse images.
The improved inference provided by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational posterior distribution, as well as sampling from the approximate posterior distribution. Table 4 reports the BER and WER for MNIST. Remarkably, for the same number of latent bits, coded models reduce both the BER and WER w.r.t. the uncoded case. Note also that the error rates grow with the number of latent bits, but this is expected due to the increased complexity of the inference process.